OpenStudy (anonymous):

write as a single fraction. simplify. -8b-3y/ 8b + 5b+2y/ 2b +1

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

Is this it? \(\dfrac{-8b-3y}{8b} + \dfrac{5b+2y}{2b+1}\)

OpenStudy (anonymous):

the negative in front of the 8b is in front of the entire equation and the +1 is seperate

OpenStudy (anonymous):

Oh, so it's this: \(-\dfrac{8b-3y}{8b} + \dfrac{5b + 2y}{2b} + 1\)

OpenStudy (anonymous):

yes =)

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Ohh this is a toughiee

OpenStudy (luigi0210):

Okay, so to do this you have to get common denominators. Do you know how to do that?

OpenStudy (anonymous):

would it be four?

OpenStudy (anonymous):

oh wait I accidentally posted this twice hold on

OpenStudy (luigi0210):

Actually the common factors would be 8b, so just multiply the second by 4 to get 8b, and the 1 will be 8b/8b, like so: \(\LARGE -\frac{8b-3y}{8b} +\frac{5b+2y}{2b} (4) + \frac{8b}{8b} \) Simplify: \(\LARGE -\frac{8b-3y}{8b} +\frac{20b+8y}{8b} + \frac{8b}{8b} \) Now get it all over a common denominator like so: \(\LARGE \frac{(20b-8b)+(8y+3y)+(8b)}{8b}\) I also distributed the negative out So it'll just be \(\LARGE \frac{20b+11y}{8b}\) Make sense?

OpenStudy (anonymous):

yes =)